OpenStudy (anonymous):

decide whether or not the equation below has a circle as its graph. 49x^2+49y^2-56x+56y-49

Parth (parthkohli):

do you know the equation of a circle?

OpenStudy (anonymous):

(x-h)^2 + (y-k)^2 = r^2

Parth (parthkohli):

There's another form: the general one.\[\rm x^2 + y^2 +Cx + Dy + E = 0\]

Parth (parthkohli):

Is the equation in your question in general form?

OpenStudy (anonymous):

yeah

Parth (parthkohli):

Well, do a thing... let's divide each term by 7.

Parth (parthkohli):

\[\rm 49x^2 -56x + 49y^2 - 56y = 49\]so,\[\rm 7x^2 - 8x + 7y^2 - 8y = 7\]factoring,\[\rm x(7x - 8) + y(7y - 8) = 7\]Hmm we already know that this ain't a circle